2. A volume of 34.7 mL of H2O is initially at 28.0 oC. A chilled glass marble weighing 4.00 g with a heat capacity of 3.52 J/oC is placed in the water. If the final temperature of the system is 26  oC , what was the initial temperature of the marble? Water has a density of 1.00 g/mL and a specific heat of 4.18 J/goC.

Enter your answer numerically, to three significant figures and in terms of oC.
Answer #1

m(marble) = 4.0 g

T(marble) = to be calculated

Since density of water is 1.00 g/mL and volume is 34.7 mL,

m(water) = 34.7 g

T(water) = 28.0 oC

C(water) = 4.18 J/goC

T = 26.0 oC

We will be using heat conservation equation

use:

heat gained by water = heat lost by marble

m(water)*C(water)*(T-T(water)) = m(marble)*C(marble)*(T(marble)-T)

34.7*4.18*(28.0-26.0) = 4.0*3.52*(26.0-T(marble))

T(marble)= 5.40 oC

Answer: 5.40 oC